Integrate sec^xlog(sinx)dx
1
Expert's answer
2020-09-29T17:58:58-0400

"\int sec^x[ln (sin x)]dx = \int sec^{x-2} [ln (sin x)] sec^{2} [ln (sin x)] dx"


Now, we apply integration by Parts by the formula;


"\int udv = uv - \int vdu"


"u = sec^{x-2} [ln (sin x)]"

"dv =sec^{2} [ln (sin x)] dx"

"du = (x-2) sec^{x-3} [ln (sin x)] . sec [ln (sin x)] .tan [ln (sin x)]"

"v = tan [ln (sin x)]"


"\therefore" "\int sec^x[ln (sin x)]dx = sec^{x-2} [ln (sin x)] tan [ln (sin x)] - (x-2) \int sec^{x-2} [ln (sin x)] tan^2 [ln (sin x)] dx"

But, "tan^2x = sec^2x -1"


"\int sec^x [ln (sin x)]dx = sec^{x-2} [ln (sin x)] tan [ln (sin x)] - (x-2) \int sec^x [ln (sin x)]dx + (x-2) \int sec^{x-2} [ln (sin x)] dx"


"\int sec^x [ln (sin x)]dx + (x-2) \int sec^x [ln (sin x)]dx = sec^{x-2} [ln (sin x)] tan [ln (sin x)] - \sout {(x-2) \int sec^x [ln (sin x)]dx} + (x-2) \int sec^{x-2} [ln (sin x)] dx + \sout{(x-2) \int sec^x [ln (sin x)]dx}"

"\frac{\sout{(x-1)} \int sec^x [ln (sin x)]dx}{\sout{(x-1)}} = \frac{ sec^{x-2} [ln (sin x)] tan [ln (sin x)]}{x-1} + \frac{x-2 \int sec^{x-2} [ln (sin x)] dx}{x-1}"


"Therefore,"

"\int sec^x [ln (sin x)]dx = \frac{ sec^{x-2} [ln (sin x)] tan [ln (sin x)]}{x-1} + \frac{x-2 }{x-1} \int sec^{x-2} [ln (sin x)] dx"

"Where, x \not = 1"